# Retinex-RAWMamba: Bridging Demosaicing and Denoising for Low-Light RAW Image Enhancement

### v4

@@ 3.4 Retinex Decomposition
Figure 4: The visualization results between our method and the state-of-the-art methods (Zoom-in for best view).
3.5 Dual-domain Encoding Stage Enhance Branch
@@
3.6 Domain Adaptive Fusion
The details of DAF are shown in Fig. 2 (c), previous feature map will be firstly concatenated with current feature map at the same level, and this result will be multiplied with previous feature map after the convolution, then it will pass through a convolution with a residual addition. And we can get the fused feature map after a final convolution. Specifically, for the two feature maps $\textbf{{pre}}\in{\mathbb{R}}^{H\times W\times C}$ and $\textbf{{cur}}\in{\mathbb{R}}^{H\times W\times C}$, they will be fused as following:
@@ 3.6 Domain Adaptive Fusion
3.7 Loss Function
Traditional low-level vision tasks generally use L1 Loss, and we also follow that, but our task involves different sub-tasks on two domains, Raw domain and sRGB domain, so the loss can be expressed as following:
@@ 3.7 Loss Function
Table 1: Quantitative results of RAW-based LLIE methods on the Sony and Fuji subsets of SID. The top-performing result is highlighted in red, while the second-best is shown in green. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better. ‘-’ indicates the result is not available.
Category Method Venue #Params.(M) Sony Fuji
@@
Single-Stage SID CVPR2018 7.7 28.96 0.787 0.356 26.66 0.709 0.432
DID ICME2019 2.5 29.16 0.785 0.368 - - -
SGN ICCV2019 19.2 29.28 0.790 0.370 27.41 0.720 0.430
LLPackNet BMVC2020 1.2 27.83 0.755 0.541 - - -
RRT CVPR2021 0.8 28.66 0.790 0.397 26.94 0.712 0.446
Multi-Stage EEMEFN AAAI2020 40.7 29.60 0.795 0.350 27.38 0.723 0.414
LDC CVPR2020 8.6 29.56 0.799 0.359 27.18 0.703 0.446
MCR CVPR2022 15.0 29.65 0.797 0.348 - - -
RRENet TIP2022 15.5 29.17 0.792 0.360 27.29 0.720 0.421
DNF CVPR2023 2.8 30.62 0.797 0.343 28.71 0.726 0.391
Ours - 6.2 30.76 0.810 0.328 29.02 0.743 0.382
4 Experiments
4.1 Datasets and Experiments Environments
4.1.1 SID Dataset
@@ 4.1.2 MCR Dataset
4.1.3 Implementation Details
@@
During training, the batch size is 1 and the initial learning rate is 1e-4, and we use the cosine annealing strategy to reduce it to 1e-5 at the 200th epoch. Adamw optimizer is used and the betas parameter is [0.9,0.999] and the momentum is 0.9. The training and testing is completed by a NVIDIA 3090 (24G), A40 (48G),respectively due to the limitation of GPU memory. We also provide the code of merging test on a 24G GPU. Note that the results of merging test are a litter bit smaller than that testing with whole image. And we use PSNR, SSIM [36] and LPIPS [47] as the quantitative evaluation metrics.
4.2 Comparison with State-of-the-Arts
We conduct experiments on SID [5] dataset including Sony and Fuji subsets and MCR [7] dataset, and compare with previous SOTA methods including SID [5], DID [26], SGN [12], EEMEFN [50], LDC [42], LLPackNet [20], RRT [19], MCR [7], RRENet [15] and DNF [17].
The results are presented in Tab. 1 and 2. As observed, most single-stage methods underperform compared to multi-stage methods, demonstrating the feasibility and effectiveness of the multi-stage approach for noisy RAW to clean sRGB cross-domain mapping. On the SID dataset, our proposed method outperforms all metrics among multi-stage approaches, while maintaining a smaller parameter count. Specifically, on the Sony and Fuji subsets, our method achieves a PSNR increase of 0.14 dB and 0.31 dB, respectively, an SSIM improvement of 0.011 and 0.017, and an LPIPS reduction of 0.015 and 0.009, compared to the best existing method.
For the MCR dataset, as shown in Tab. 2, while our improvement in SSIM is modest, we achieve a significant PSNR increase of 1.14 dB, a 3.6% enhancement over the second-best method. Additionally, we selected several previous state-of-the-art (SOTA) methods and visualized their performance on the SID Sony dataset, as shown in Fig. 4. Three scenarios are depicted, each containing two sub-regions. In the first two scenes, most other methods produce a green tint to the image. In the third scene, these methods often fail to preserve details adequately. In contrast, our proposed method closely aligns with the ground truth in both color and detail, effectively achieving denoising and color enhancement in the raw domain under low-light conditions.
4.3 Ablation Studies
@@
Table 2: Quantitative results on MCR [7] dataset. The top-performing result is highlighted in red, while the second-best is shown in green. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better.
Category Method PSNR $\uparrow$ SSIM $\uparrow$
Single-Stage SID 29.00 0.906
DID 26.16 0.888
SGN 26.29 0.882
RRT 25.74 0.851
Multi-Stage LDC 29.36 0.904
MCR 31.69 0.908
DNF 32.00 0.915
Ours 33.14 0.914
To demonstrate the validity of our proposed method, we perform ablation experiments on the SID Sony dataset. We first propose a baseline model that consists only of SDB and the unmodified naive visual mamba in MambaIR [13] and GFM in DNF [17]. Tab. 3 shows the results of adding or replacing the corresponding module based on the baseline, where RAWM stands for replacing naive Mamba with RAWMamba, RDM stands for adding RDM module, and DAF stands for replacing GFM with DAF module. All the ablation experiments were conducted in the same environment.
First, we replaced the baseline’s naive Mamba with the proposed RAWMamba. The results showed increases of 0.41 dB in PSNR and 0.012 in SSIM, demonstrating that our RAWMamba, with its eight-directional scanning mechanism, performs well in the demosaicing task. Next, we incorporated the proposed RDM for denoising and automatic exposure correction. The results indicated that although SSIM did not improve, PSNR increased by an additional 0.27 dB. This suggests that the initial exposure of the images was indeed problematic, and our RDM effectively enhances denoising and exposure correction. Finally, we replaced all GFM components in the network with our proposed DAF to improve the stability of the training process. This led to further gains, with PSNR and SSIM increasing by 0.06 dB and 0.001, respectively.
Furthermore, we conducted a straightforward visualization of the ablation study, depicted in Fig. 5. The incorporation of RAWMamba into the baseline model effectively reduces noise and mitigates the green color distortion. Comparatively, the Retinex-RawMamba approach demonstrates superior color correction capabilities and attains the highest PSNR and SSIM scores. This clearly indicates that our proposed method outperforms others in terms of both detail preservation and color accuracy.
Table 3: Ablation study on SID Sony dataset.
Baseline RAWM RDM DAF PSNR SSIM
30.04 0.797
30.45 0.809
30.70 0.809
30.76 0.810
Figure 5: The visualization results for ablation studies (Zoom-in for best view).
5 Conclusion
For the task of denoising and enhancing RAW images under low-light conditions, we introduce Retinex-RAWMamba, a novel two-stage cross-domain network. Our approach extends the capabilities of the traditional Vision Mamba by incorporating RAWMamba, which exploits the inherent properties of demosaicing algorithms in ISP to achieve enhanced color correction and detail retention. Additionally, we integrate Retinex theory through our Retinex Decomposition Module, facilitating automatic exposure correction and yielding RGB images with improved illumination and brightness fidelity. Comprehensive theoretical analysis and experimental validation underscore the effectiveness and significant potential of our method.

### v5

@@ III-D Retinex Decomposition and Dual-domain Encoding Stage Enhance Branch
Figure 4: The visualization results between our method and the state-of-the-art methods, and the red and green box areas are cropped out for easy comparison (Zoom-in for best view).
@@
III-E Domain Adaptive Fusion
The details of DAF are shown in Fig. 2 (c), previous feature map will be firstly concatenated with current feature map at the same level, and this result will be multiplied with previous feature map after the convolution, then it will pass through a convolution with a residual addition. And we can get the fused feature map after a final convolution. Specifically, for the two feature maps $\textbf{{pre}}\in{\mathbb{R}}^{H\times W\times C}$ and $\textbf{{cur}}\in{\mathbb{R}}^{H\times W\times C}$, they will be fused as follows:
@@ III-E Domain Adaptive Fusion
III-F Loss Function
Traditional low-level vision tasks generally use L1 Loss, and we also follow that, but our task involves different sub-tasks on two domains, Raw domain and sRGB domain, so we use L1 loss for each domain to better guide model learning. And the total loss can be expressed as follows:
@@ III-F Loss Function
TABLE I: Quantitative results of RAW-based LLIE methods on the Sony and Fuji subsets of SID. The top-performing result is highlighted in bold, while the second-best is shown in underline. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better. ‘-’ indicates the result is not available.
Category Method Venue #Params.(M) GFLOPs Sony Fuji
@@
Single-Stage SID[9] CVPR2018 7.7 48.5 28.96 0.787 0.356 26.66 0.709 0.432
DID[1] ICME2019 2.5 669.2 29.16 0.785 0.368 - - -
SGN[2] ICCV2019 19.2 75.5 29.28 0.790 0.370 27.41 0.720 0.430
LLPackNet[67] BMVC2020 1.2 7.2 27.83 0.755 0.541 - - -
RRT[3] CVPR2021 0.8 5.2 28.66 0.790 0.397 26.94 0.712 0.446
Multi-Stage EEMEFN[31] AAAI2020 40.7 715.6 29.60 0.795 0.350 27.38 0.723 0.414
LDC[4] CVPR2020 8.6 124.1 29.56 0.799 0.359 27.18 0.703 0.446
MCR[5] CVPR2022 15.0 90.5 29.65 0.797 0.348 - - -
RRENet[30] TIP2022 15.5 96.8 29.17 0.792 0.360 27.29 0.720 0.421
Ma et al.[68] NN2023 - - 29.38 0.793 0.387 27.40 0.722 0.505
DNF[10] CVPR2023 2.8 57.0 30.62 0.797 0.343 28.71 0.726 0.391
Ours - 6.2 113.6 30.76 0.810 0.328 29.02 0.743 0.382
IV Experiments
IV-A Datasets and Experiments Environments
IV-A1 SID Dataset
@@ IV-A2 MCR Dataset
IV-A3 ELD Dataset
The ELD[27] dataset contains 10 indoor scenes and 4 camera devices from multiple brands (i.e., SonyA7S2, NikonD850, CanonEOS70D, CanonEOS700D). We choose the commonly used SonyA7S2 and NikonD850 with three ISO levels (800, 1600, and 3200)5 and two low light factors (100, 200) for validation, resulting in 120 (3×2×10×2) raw image pairs in total. And we chose SID, MCR and DNF for comparison, using the pretrained model on Sony-SID dataset to compare the generalization.
IV-A4 LOL Dataset
The LOL dataset has v1[69] and v2[70] versions. LOL-v2 is divided into real and synthetic subsets. The training and testing sets are split in proportion to 485:15, 689:100, and 900:100 on LOL-v1, LOL-v2-real, and LOL-v2-synthetic.
IV-A5 Implementation Details
@@
During training, the batch size is 1 and the initial learning rate is 1e-4, and we use the cosine annealing strategy to reduce it to 1e-5 at the 200th epoch. The total number of epochs is 250. Adamw optimizer is used and the betas parameter is [0.9,0.999] and the momentum is 0.9. The training and testing is completed by a NVIDIA 3090 (24G), A40 (48G),respectively due to the limitation of GPU memory. We also provide the code of merging test on a 24G GPU. Note that the results of merging test are a litter bit smaller than that testing with whole image. And we use PSNR, SSIM [71] and LPIPS [72] as the quantitative evaluation metrics.
IV-B Comparison with State-of-the-Arts
@@
TABLE II: Quantitative results on MCR [5] dataset. The top-performing result is highlighted in bold, while the second-best is shown in underline. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better. The inference time is computed for the entire MCR test set.
Category Method PSNR $\uparrow$ SSIM $\uparrow$ Time(s) $\downarrow$
Single-Stage SID [9] 29.00 0.906 35
DID [1] 26.16 0.888
SGN [2] 26.29 0.882
RRT [3] 25.74 0.851
Multi-Stage LDC [4] 29.36 0.904
MCR [5] 31.69 0.908 48
DNF [10] 32.00 0.915 72
Ours 33.14 0.914 147
We conduct experiments on SID [9] dataset including Sony and Fuji subsets and MCR [5] dataset, and compare with previous SOTA methods including SID [9], DID [1], SGN [2], EEMEFN [31], LDC [4], LLPackNet [67], RRT [3], MCR [5], RRENet [30], DNF [10], and Ma et al. [68].
Figure 5: Visualization results of MCR dataset, and the format of the value under each visualized image is “PSNR / SSIM” (Zoom-in for best view).
The results are presented in Tab. I and II. As observed, most single-stage methods underperform compared to multi-stage methods, demonstrating the feasibility and effectiveness of the multi-stage approach for noisy RAW to clean sRGB cross-domain mapping. Compared to the most recent work [68], we can see that for Sony dataset, our method achieve improvements of 1.38 and 0.017 on PSNR and SSIM, respectively, and 1.62 and 0.021 for Fuji dataset. Overall, on the SID dataset, our proposed method outperforms all metrics among multi-stage approaches, while maintaining a smaller parameter count. Specifically, on the Sony and Fuji subsets, our method achieves a PSNR increase of 0.14 dB and 0.31 dB, respectively, an SSIM improvement of 0.011 and 0.017, and an LPIPS reduction of 0.015 and 0.009, compared to the best existing method.
For the MCR dataset, as shown in Tab. II, while our improvement in SSIM is modest, we achieve a significant PSNR increase of 1.14 dB, a 3.6% enhancement over the second-best method. Note that we did not include direct comparisons with some existing methods due to the lack of publicly available code and the fact that those methods were not originally evaluated on the MCR dataset. As such, reproducing their results under consistent settings would be unreliable. To further support the generalization ability of our method, we conducted additional experiments on the ELD [27] and LOL [69] datasets.
As for the efficiency of the proposed model, we can see that in Tab. I and Tab. II. Specifically, 113.6 GFLOPs for $4\times 512\times 512$ input, and 147 s for the entire MCR test set. Notably, while our approach is not optimized for real-time applications, it prioritizes image quality—a critical factor in real-world scenarios like photo editing or professional imaging systems, where output fidelity often outweighs inference speed. That said, we recognize the importance of improving inference efficiency in vision Mamba architectures. As outlined in our future work, we plan to explore lightweight design adaptations to balance performance and quality in subsequent research.
Additionally, we selected several previous state-of-the-art (SOTA) methods and visualized their performance on the SID Sony dataset, as shown in Fig. 4. Three scenarios are depicted, each containing two sub-regions. In the first two scenes, most other methods produce a green tint to the image. In the third scene, these methods often fail to preserve details adequately. In contrast, our proposed method closely aligns with the ground truth in both color and detail, effectively achieving denoising and color enhancement in the raw domain under low-light conditions. More visualization results are shown in Fig. 6 and 7. Moreover, a few visualization results of MCR dataset are also presented in Fig. 5.Taking the last column as an example, in comparison with the ground truth (GT), the MCR method exhibits a significant deviation in color, with an overall reddish bias. The DNF results demonstrate noticeable over-exposure in the lower central region. In contrast, our proposed method achieves markedly superior performance, evidenced by substantial improvements in key metrics such as PSNR and SSIM. Note that the format of the value under each visualized image is “PSNR / SSIM”.
Moreover, to explore the generalization of our proposed method, we also conducted experiments on ELD[27] dataset. We used the pretrained models from SID[9], MCR[5], DNF[10] and ours that were trained on SID-Sony dataset, and obtained the results on ELD dataset, which is shown in Tab. IV. For SonyA7S2 dataset, our pretrained model achieves highest PSNR and comparable SSIM with DNF. Specifically, 0.54 and 1.27 improvements in PSNR for ratio 100 and 200, respectively. As for NikonD850 dataset, our model achieves highest SSIM for both ratio. For ratio 100, we get comparable PSNR with DNF, and for ratio 200, we get 1.13 improvement in PSNR. Note that the Raw images in SID-Sony and ELD-SonyA7S2 were captured with the same model camera. According to the experiments’ results, we can conclude that our method has better generalization especially for different cameras input images than other end-to-end models.
TABLE III: Quantitative results on LOL datasets. The top-performing result is highlighted in bold, while the second-best is shown in underline.
Methods Complexity LOL-v1 LOL-v2-real LOL-v2-syn
FLOPs (G) Params (M) PSNR SSIM PSNR SSIM PSNR SSIM
SID [9] 13.73 7.76 14.35 0.436 13.24 0.442 15.04 0.610
UFormer [73] 12.00 5.29 16.36 0.771 18.82 0.771 19.66 0.871
RetinexNet [69] 587.47 0.84 16.77 0.560 15.47 0.567 17.13 0.798
Chen et al. [74] 18.31 0.489
RetinexNet+LPDM [75] 18.03 0.760
Restormer [76] 144.25 26.13 22.43 0.823 19.94 0.827 21.41 0.830
MIRNet [77] 785.00 31.76 24.14 0.830 20.02 0.820 21.94 0.876
Proposed Method (Two-Stage) 103.76 6.22 23.23 0.805 21.14 0.779 24.93 0.924
Proposed Method (Second Stage Only) 54.54 2.87 23.07 0.807 20.91 0.800 25.36 0.929
TABLE IV: Quantitative results on ELD [27] dataset. The top-performing result is highlighted in bold, while the second-best is shown in underline. Higher ratio represents more noise.
Dataset Ratio SID [9] MCR[5] DNF[10] Ours
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
SonyA7S2 100 26.80/0.826 26.62/0.837 28.35/0.869 28.89/0.861
200 25.98/0.780 24.00/0.717 26.94/0.819 28.21/0.818
Avg 26.39/0.803 25.31/0.777 27.64/0.844 28.55/0.840
NikonD850 100 25.77/0.764 25.60/0.785 27.51/0.816 27.47/0.838
200 25.01/0.739 25.41/0.758 25.93/0.769 27.06/0.806
Avg 25.39/0.751 25.50/0.772 26.72/0.792 27.27/0.822
Additionally, to validate the models’ generalization to low light RGB image, we additionally conducted experiments on LOL datasets. The results are shown in Tab. III. Since the proposed method is not for RGB inputs originally, the training strategy is different. We conducted two experiments with different settings for all three LOL datasets. Specifically, “Two-Stage” just used the original two-stages model without the intermediate supervision, while “Second Stage Only” removed the first stage and just used the second stage, leading to less GFLOPs and parameters. We can see that for all three dataset, we achieved comparable performance with some previous method. It’s worth noting that “Second Stage Only” even performs better than “Two-Stage” with less parameters, which indicates that the effectiveness of first stage is weakened by removing the supervision. Although we achieved lower PSNR and SSIM than MIRNet on LOL-v1 and LOL-v2-real dataset, our “Second Stage Only” achieve much higher PSNR and SSIM on LOL-v2-sys dataset, which improves by 3.42 and 0.053, respectively, with 6.9% GFLOPs and 9% parameters of MIRNet. However, we still want to emphasize that our method is specialized for low light RAW images. How to better apply it to RGB images is worth further research in the future.
[…]
Figure 6: More visualization results (Zoom-in for best view).
TABLE V: Feature analysis of LOL dataset. The greater the number of layers, the deeper the layers. The results are presented as the respective cosine similarity of [L,E]/[H,E]. B, O represents “Baseline” and “Ours”, respectively.
Dataset Layers 1 2 3 Avg
V1 B 0.6234/0.8111 0.9827/0.9897 0.9701/0.9803 0.8587/0.9270
O 0.6518/0.8344 0.9970/0.9989 0.9847/0.9971 0.8778/0.9435
V2-R B 0.6164/0.8033 0.9753/0.9845 0.9774/0.9811 0.8564/0.9230
O 0.6190/0.8054 0.9970/0.9991 0.9875/0.9978 0.8678/0.9341